Bobo, Thabo and Khaya are paid R240 for moving the lawn
The split the money in ratio. 5:9.10
1. How much less did Bobo make than Thabo?
2. Bobo complains about making so much less. The boys decides to divide their pay in a ratio of 3:4:5. Howuch more does Bobo make than (1)?
1
Expert's answer
2021-07-19T13:42:26-0400

bobo , than (thabo) and khanya ratio for 240 rs is

5:9:10

let bobo be 5x

than(thabo) be 9x

and khanya be 10x

so , 5x+9x+10x = 240

24x = 240

x = 10

so bobo = 5 (10) = 50 rs

than(thabo) = 9(10) = 90 rs

khanya= 10 (10) = 100 rs

1.1 "\implies"

how much less did Bobo make than thabo?

thabo price = 90 rs and bobo price = 50 rs

= 90 - 50 = 40 rs


1.2 "\implies"

ratio of b0bo , than(thabo) and khanya for 240 rs is changed to

3:4:5

let be bobo = 3x

thabo = 4x

khanya = 5x

so ,

3x + 4x + 5x = 240 rs

12x = 240

x = 20

than we got , bobo = 3(20) = 60 rs

thabo = 4(20) = 80rs

khanya = 5(20) = 100rs

so according to 1.1 bobo gives 50 rs and this part bobo gives 60 rs

so , this part price - previous part price

= 60 - 50 = 10 rs

10rs more give than 1.1